can you help me find the roots of the quadratic equation x2+6x-41=0 and walk me through step by step
Answer by lwsshak3(11628)

find the roots of the quadratic equation
x^2+6x-41=0
complete the square:
(x^2+6x+9)-9-41=0
(x+3)^2=50
x+3=±√50
x=-3±√50
x=-3-√50
or
x=-3+√50